# Scalable non-separable spatio-temporal Gaussian process models for large-scale short-term weather prediction

### v1

@@ Appendix F Real-world application: Additional plots and tables
Table 7: Overall five-day RMSE, CRPS, and average computational runtimes. Training time corresponds to parameter estimation; prediction time corresponds to prediction of means and variances.
@@ F.1 PIT reliability diagrams (temperature)
Table 9: The average overall MAE, CRPS, BS, and LS for the Vecchia approximation with Euclidean- and correlation-based neighbor selection, the FITC approximation with inducing points selected via standard and space-time separated kMeans++, and the VIF approximation. The results distinguish between temporal forecasting, and spatio-temporal forecasting. Average runtimes in seconds (s) are shown for parameter estimation (Train) and computation of predictive mean and variances (Prediction).

### v2

@@ Appendix F Real-world application: Additional plots and tables
Table 7: Average five-day RMSE, CRPS, and average computational runtimes for temperature modeling. Training time corresponds to parameter estimation; prediction time corresponds to prediction of means and variances. The training datasets contain between approximately $800,000$ and $1.3$ million space-time points.
@@ F.1 PIT reliability diagrams (temperature)
Table 9: Average five-day MAE, CRPS, BS, and LS, and average computational runtimes for precipitation modeling. The results distinguish between temporal forecasting and spatio-temporal forecasting. Average runtimes in seconds (s) are shown for parameter estimation (Train) and computation of predictive mean and variances (Prediction). The training datasets contain between approximately $800,000$ and $1.3$ million space-time points.